Please explain this linear equation:
If x + y = 15 and x -y = -19 then xy = ?
Thanks,
Jesse Evans
jesseevans@hughes.net

Found 2 solutions by papugai, Alan3354:
Answer by papugai(18)

If x + y = 15 and x -y = -19 then xy = ?
Add the two equations:
2x=-4
x=-2
sub into first equation:
-2+y=15
y=17
xy=-2*17=-34
Hope this was helpful

Answer by Alan3354(69443)

The other tutor solved it, but just a note:
it's not a good idea to show your complete name and your email.